Find out cost of Goods sold and Sale price per unit in Rs. from the following information. Show your computations mathematically.

Opening stock ?

Average inventory 1, 80,000

Direct Expenses 150% of closing stock

Purchases 3, 20,000

Returns outwards 40,000

Units sold 1010

Closing stock is three times the opening stock.

Gross profit margins are 16.67% on sales value.


1
Expert's answer
2021-11-01T19:15:36-0400

Calculating the opening stock and the closing stock:Average inventory = 180000

(Opening stock + Closing stock )/ 2 = 180000

(Opening stock + 3*Opening stock) / 2 = 180000

Opening stock = 90000

Closing stock = 270000


Calculating the Cost of goods sold:Cost of goods sold = Opening stock + Purchases - Returns outwards - Closing stock

                = 90000 + 320000 - 40000 - 270000

                = 100000


Calculating the Sale value:Sales - Sales * 16.67% = Cost of goods sold

Sales - Sales * 16.67% = 100000

Sales = 120000


Calculating the Sale price per unit:Sale price per unit = Sales value / units sold

               = 120000 / 1010

               = Rs 118.81